# Dissociation of Faithful and Unfaithful Reasoning in LLMs

### v1

@@ 8 Discussion
In this study, we have investigated the ability of LLMs to recover from errors in their reasoning. We challenge the assumption that error recovery indicates unfaithful reasoning per-se, and find that many instances of error recovery are fully justified and interpretable within the CoT transcript. We develop a more fine-grained annotation scheme and analysis of error recovery than has been performed by other studies, which is supported by extensive manual annotation of LLM reasoning transcripts.
We identify three key aspects of chain of thought transcripts which seem to have an effect on these language models’ ability to recover. First, Experiment 1 finds that evidence for the correct value in the prior context increases the error recovery rate. Furthermore, experiment 2 finds that larger errors are easier to recover from. Finally, experiment 3 finds that stronger prior expectations that an error will occur increase the frequency of recovery.
The experiments also evaluated the effect of these interventions on faithful and unfaithful error recoveries. In all 3 experiments, the factors that increase overall recoveries also increase faithful recoveries, but not unfaithful recoveries.
@@ 9 Conclusion
A major goal in LLM interpretability research is to understand how these systems reason. If chain of thought transcripts faithfully reflect the internal reasoning process of models, that would greatly increase our ability to monitor these models and control them.
The current paper finds that it is easy to induce unfaithful reasoning in LLMs by forcing error recovery. Significantly, the factors that increase the rate of faithful error recoveries were found to decrease the rate of unfaithful error recoveries. This provides evidence for dissociation between faithful and unfaithful reasoning.
In psychological studies, dissociation provides evidence for distinct mechanisms for cognitive processes. For example, Tulving (1972) identified experimental tasks which selectively interfere with episodic memory or semantic memory, but not both. This differential response to these experimental tasks suggested that these are distinct types of memory, with different underlying cognitive processes.
In the current study, the dissociation between faithful and unfaithful recoveries provides evidence for multiple reasoning mechanisms. Faithful recoveries occur more frequently when the model has strong evidence that an error occurred. Unfaithful recoveries occur more often when there is less evidence of an error. These findings leave open the precise mechanisms underlying faithful and unfaithful reasoning, and how these mechanisms arose during training.
@@ 10 Limitations
The study is limited in its scale. We evaluated two models from a single model class (GPT), and in the main body of the work, we provide analysis for only 1 of those models (GPT-4). We chose to focus on depth over breadth, leveraging costly manual annotations and providing specific analysis for each case.
Additionally, although we were able to achieve a high level of control over our perturbations in these experiments, we cannot be sure that the errors we introduced are representative of errors which these models would naturally produce. If our errors are too far out of distribution for the models, our findings may not apply to in-distribution errors.
On a more conceptual level, this work shares a broad limitation with many faithfulness investigations; we do not have access to these models’ weights and activations, or a reliable way to directly interpret these representations.
@@ 11 Reproducibility
Code and data for our experiments, including instructions for reproducing our results, will be made available at https://github.com/CoTErrorRecovery/CoTErrorRecovery. However, OpenAI has announced that access to the GPT-3.5 and GPT-4 checkpoints that we evaluated (gpt-3.5-turbo-0301 and gpt-4-0314) may be permanently deprecated as early as June 2024 (Eleti et al., 2023).
@@ A.4.5 Annotators
Appendix B Full error recovery results
B.1 Experiment 1 full results
@@
Figure 10: Total error recovery, faithful recovery, and unfaithful recovery rates for GPT-3.5 in experiment 1. All values are shown as percentage of total validated responses per condition. Error bars indicate 95% binomial confidence intervals.
@@
Table 2: Experiment 1 numerical results. All percentages are rounded to 2 decimal places.
Dataset Error
@@
MultiArith Copy Error GPT-3.5 288 77.78 6.60 67.71 3.47
GPT-4 295 99.32 72.54 22.37 4.41
Calculation Error GPT-3.5 264 22.35 14.39 7.58 0.38
GPT-4 247 77.73 68.42 4.45 4.88
Propagated Calculation
Error
GPT-3.5 280 15.00 13.93 0.71 0.36
GPT-4 295 32.54 25.42 6.44 0.68
ASDiv Copy Error GPT-3.5 287 78.05 10.10 65.85 2.09
GPT-4 287 97.21 66.90 20.91 9.66
Calculation Error GPT-3.5 225 23.56 3.11 18.67 1.78
GPT-4 229 74.67 51.97 18.78 3.93
Propagated Calculation
Error
GPT-3.5 226 12.39 5.31 7.08 0.00
GPT-4 204 52.94 16.67 35.78 0.49
SVAMP Copy Error GPT-3.5 284 70.77 6.34 63.73 0.70
GPT-4 284 97.89 66.90 21.13 10.07
Calculation Error GPT-3.5 246 16.67 4.47 10.57 1.63
GPT-4 266 78.57 56.39 20.30 1.89
Propagated Calculation
Error
GPT-3.5 273 5.49 2.20 3.30 0.00
GPT-4 254 53.94 13.78 40.16 0.00
GSM8K Copy Error GPT-3.5 293 64.51 9.90 47.10 7.51
GPT-4 284 95.07 58.10 20.77 15.81
Calculation Error GPT-3.5 266 10.15 6.01 3.01 1.13
GPT-4 227 78.85 52.42 9.69 16.38
Propagated Calculation
Error
GPT-3.5 292 5.82 5.14 0.34 0.34
GPT-4 247 26.72 12.55 13.77 0.38
B.2 Experiment 2 full results
Figure 11: Overall error recovery rates for GPT-4 with small errors and large errors. Error bars indicate 95% binomial confidence intervals.
Figure 12: Faithful recovery rates for GPT-4 with small errors and large errors. Error bars indicate 95% binomial confidence intervals.
Figure 13: Unfaithful recovery rates for GPT-4 with small errors and large errors. Error bars indicate 95% binomial confidence intervals.
Table 3: Experiment 2 numerical results, performed on GPT-4. All percentages are rounded to 2 decimal places.
Dataset Error Position Error Amount n Total Error Recovery (%) Faithful Recovery (%) Unfaithful Recovery (%)
Complete Hallucination (%) Partial Hallucination (%)
MultiArith Copy Error Small Error 288 97.57 63.19 30.21 4.17
Large Error 288 99.65 90.62 7.64 1.39
Calculation Error Small Error 247 77.33 72.06 3.64 1.62
Large Error 247 100.00 95.95 0.81 3.24
Propagated Calculation Error Small Error 295 32.88 26.10 6.44 0.34
Large Error 295 95.25 83.05 12.20 0.00
ASDiv Copy Error Small Error 290 97.59 61.38 29.31 6.90
Large Error 290 98.62 85.17 10.69 2.76
Calculation Error Small Error 229 75.55 47.16 24.45 3.93
Large Error 229 97.38 85.59 8.30 3.49
Propagated Calculation Error Small Error 215 49.30 12.09 37.21 0.47
Large Error 215 91.63 72.09 19.53 0.00
SVAMP Copy Error Small Error 289 98.27 66.78 21.11 10.38
Large Error 289 98.96 82.35 13.49 3.11
Calculation Error Small Error 280 75.00 52.86 20.36 1.79
Large Error 280 94.29 85.71 7.14 1.43
Propagated Calculation Error Small Error 254 39.76 4.33 35.43 0.00
Large Error 254 90.55 69.29 21.26 0.00
GSM8K Copy Error Small Error 291 96.98 49.23 29.29 18.46
Large Error 291 96.98 68.12 15.10 13.76
Calculation Error Small Error 232 79.22 51.08 6.06 22.08
Large Error 232 97.84 80.53 6.49 10.82
Propagated Calculation Error Small Error 263 24.08 12.71 11.37 0.00
Large Error 263 64.88 42.15 22.07 0.66
B.3 Experiment 3 full results
Figure 14: Overall error recovery rates for GPT-4 with textual adjustments. The ”Baseline” results are the same error recovery results from Experiment 1 (Figure 4). Error bars indicate 95% binomial confidence intervals.
Figure 15: Faithful recovery rates for GPT-4 with textual adjustments. The ”Baseline” results are the same error recovery results from Experiment 1 (Figure 5). Error bars indicate 95% binomial confidence intervals.
Figure 16: Unfaithful recovery rates for GPT-4 with textual adjustments. The ”Baseline” results are the same error recovery results from Experiment 1 (Figure 5). Error bars indicate 95% binomial confidence intervals.
Figure 17: Difference in GPT-4’s recovery rates between error recovery prompt and baseline conditions. The baseline results are the same error recovery results from Experiment 1 (Figure 5), which had no textual adjustments. Negative values indicate recoveries occurred more often in the baseline condition. Error bars are 95% confidence intervals.
Figure 18: Overall error recovery rates for GPT-3.5 with textual adjustments. The ”Baseline” results are the same error recovery results from Experiment 1 (Figure 10). Error bars indicate 95% binomial confidence intervals.
Table 4: Experiment 3 numerical results for the MultiArith and ASDiv datasets. All percentages are rounded to 2 decimal places. Due to manual annotation costs, faithfulness analysis was not performed for the GPT-3.5 text adjustment conditions. (For ease of comparison, ”Baseline” results are copied from Table 2 in Experiment 1.)
Dataset Error Position Text Adjustment Model n Total Error Recovery (%) Faithful Recovery (%) Unfaithful Recovery (%)
Complete Hallucination (%) Partial Hallucination (%)
MultiArith Copy Error Baseline GPT-3.5 288 77.78 6.60 67.71 3.47
GPT-4 295 99.32 72.54 22.37 4.41
Context Noise GPT-3.5 281 79.00
GPT-4 295 99.32 80.68 18.31 0.33
Error Recovery Prompt GPT-3.5 288 71.18
GPT-4 295 98.64 68.81 27.80 2.03
Calculation Error Baseline GPT-3.5 264 22.35 14.39 7.58 0.38
GPT-4 247 77.73 68.42 4.45 4.86
Context Noise GPT-3.5 264 24.62
GPT-4 246 87.45 83.00 3.64 0.81
Error Recovery Prompt GPT-3.5 264 24.24
GPT-4 247 81.78 75.71 4.45 1.62
Propagated
Calculation Error
Baseline GPT-3.5 280 15.00 13.93 0.71 0.36
GPT-4 295 32.54 25.42 6.44 0.68
Context Noise GPT-3.5 280 15.71
GPT-4 295 52.20 45.08 6.44 0.68
Error Recovery Prompt GPT-3.5 280 15.71
GPT-4 295 39.66 31.53 8.14 0.00
ASDiv Copy Error Baseline GPT-3.5 287 78.05 10.10 65.85 2.09
GPT-4 287 97.21 66.90 20.91 9.40
Context Noise GPT-3.5 279 80.65
GPT-4 279 97.85 76.70 20.79 0.36
Error Recovery Prompt GPT-3.5 287 77.70
GPT-4 287 96.52 67.60 25.44 3.48
Calculation Error Baseline GPT-3.5 225 23.56 3.11 18.67 1.78
GPT-4 229 74.67 51.97 18.78 3.93
Context Noise GPT-3.5 224 28.13
GPT-4 229 87.34 64.19 23.14 0.00
Error Recovery Prompt GPT-3.5 225 27.11
GPT-4 229 81.22 55.90 24.02 1.31
Propagated
Calculation Error
Baseline GPT-3.5 226 12.39 5.31 7.08 0.00
GPT-4 204 52.91 16.67 35.78 0.49
Context Noise GPT-3.5 226 11.50
GPT-4 204 71.57 38.73 32.84 0.00
Error Recovery Prompt GPT-3.5 226 11.95
GPT-4 204 64.71 20.59 44.12 0.00
Table 5: Experiment 3 numerical results for the SVAMP and GSM8K datasets. All percentages are rounded to 2 decimal places. Due to manual annotation costs, faithfulness analysis was not performed for the GPT-3.5 text adjustment conditions. (For ease of comparison, ”Baseline” results are copied from Table 2 in Experiment 1.)
Dataset Error Position Text Adjustment Model n Total Error Recovery (%) Faithful Recovery (%) Unfaithful Recovery (%)
Complete Hallucination (%) Partial Hallucination (%)
SVAMP Copy Error Baseline GPT-3.5 284 70.77 6.34 63.73 0.70
GPT-4 284 97.89 66.90 21.13 9.86
Context Noise GPT-3.5 271 78.60
GPT-4 277 99.28 84.84 14.44 0.00
Error Recovery Prompt GPT-3.5 284 69.71
GPT-4 284 99.30 69.01 25.35 4.93
Calculation Error Baseline GPT-3.5 246 16.67 4.47 10.57 1.63
GPT-4 266 78.57 56.39 20.30 1.88
Context Noise GPT-3.5 246 21.14
GPT-4 264 89.77 70.45 19.32 0.00
Error Recovery Prompt GPT-3.5 246 19.51
GPT-4 266 85.71 60.15 20.30 5.26
Propagated
Calculation Error
Baseline GPT-3.5 273 5.49 2.20 3.30 0.00
GPT-4 254 53.94 13.78 40.16 0.00
Context Noise GPT-3.5 273 7.33
GPT-4 254 73.62 41.73 31.89 0.00
Error Recovery Prompt GPT-3.5 273 8.42
GPT-4 254 63.78 15.75 48.03 0.00
GSM8K Copy Error Baseline GPT-3.5 293 64.51 9.90 47.10 7.51
GPT-4 284 95.07 58.10 20.77 16.20
Context Noise GPT-3.5 290 70.69
GPT-4 284 96.48 70.07 25.00 1.41
Error Recovery Prompt GPT-3.5 293 66.10
GPT-4 284 92.96 53.47 30.28 10.21
Calculation Error Baseline GPT-3.5 266 10.15 6.01 3.01 1.13
GPT-4 227 78.85 52.42 9.69 16.74
Context Noise GPT-3.5 266 14.66
GPT-4 227 83.70 69.16 12.78 1.76
Error Recovery Prompt GPT-3.5 266 13.52
GPT-4 227 80.18 65.64 5.29 9.25
Propagated
Calculation Error
Baseline GPT-3.5 292 5.82 5.14 0.34 0.34
GPT-4 247 26.72 12.55 13.77 0.40
Context Noise GPT-3.5 292 6.85
GPT-4 247 51.42 33.20 17.81 0.41
Error Recovery Prompt GPT-3.5 292 9.93
GPT-4 247 30.77 16.19 14.57 0.00
Appendix C Sensitivity Analysis

### v2

@@ 8 Discussion
In this study, we have investigated the ability of LLMs to recover from errors in their reasoning. We developed a fine-grained annotation scheme for LLM errors, and performed extensive manual annotations of LLM errors across three models and four datasets.
We identified three key aspects of chain of thought transcripts which have an effect on these language models’ ability to recover. First, Experiment 1 found that larger errors are easier to recover from. Experiment 2 found that stronger prior expectations that an error will occur increase the frequency of recovery. Experiment 3 found that evidence for the correct value in the prior context increases the error recovery rate.
The experiments also evaluated the effect of these interventions on faithful and unfaithful error recoveries. Across the experiments, faithful and unfaithful recoveries consistently diverged in their responses to these interventions. Factors that increased the rate of faithful recoveries decreased the rate of unfaithful recoveries, and vice-versa.
@@ 9 Conclusion
Our study provides evidence that LLMs operate with two distinct modes of reasoning. In one mode, the model generates text that is optimized for human interpretability, allowing a human to understand the reasoning that supports the conclusion. In the other mode, the LLM arrives at conclusions through internal processes that are not fully captured in the generated text. The generated text in this mode may appear plausible but does not provide a valid argument for the model’s conclusions.
The second, unfaithful reasoning mode undermines our ability to reliably audit an LLM’s decision-making process. The model-generated explanations may appear plausible but fail to accurately represent the true basis for the model’s conclusions. This means that examining a model’s chain of thought output is not sufficient for verifying its conclusions.
A key challenge is to develop methods which consistently elicit the interpretable mode of reasoning from LLMs. Future research should focus on understanding the mechanisms underlying these different modes of reasoning, identifying factors that influence which mode is activated, and developing techniques to encourage faithful reasoning.
@@ 10 Limitations
The study evaluated 3 models on 4 datasets, which consist of mathematical reasoning problems. We chose to focus on depth over breadth, leveraging costly manual annotations for each case. However, the study does not investigate non-mathematical reasoning errors.
Additionally, errors that these models naturally produce during text generation may be different than those introduced in the current study. To address this, we analyzed a sample of model-generated errors from the math word problem datasets (see Appendix B), and found natural behavior similar to the errors and recovery behaviors investigated in this study. Further work is required to understand the full range of naturally occurring errors in LLMs.
@@ 11 Reproducibility
Code and data for our experiments, including instructions for reproducing our results, will be made available at https://github.com/CoTErrorRecovery/CoTErrorRecovery. However, OpenAI has announced that access to the GPT-4 checkpoint that we evaluated (gpt-4-0314) may be permanently deprecated as early as June 2024 (Eleti et al., 2023).
@@ A.4.5 Annotators
Appendix B Analysis of Natural Errors
In our experiments, we prompted LLMs to generate reasoning for solving math problems, and then introduced artificial errors into the reasoning transcripts. In this section, we analyze naturally generated LLM errors to determine whetherthey match the distribution of natural errors. We filtered these transcripts to ensure that they were valid chains of reasoning, and only introduced errors into the valid transcripts.
We collected reasoning transcripts from GPT-4, and identified transcripts that contained incorrect reasoning. [^4] These invalid transcripts were excluded from all other analyses in the paper. We then manually labeled 50 of these transcripts per dataset (200 total). After filtering out low-quality questions (e.g. where the ground-truth target answer is ambiguous or incorrect) and subjectively correct answers (e.g. where the model reached the correct answer at some point in its reasoning but did not output the answer in the correct format), we found 74 high-quality, naturally-generated incorrect chains of thought. Of these, 7 transcripts involved the introduction and propagation of calculation errors, as defined in our paper (see Section 7). Additionally, 6 more error transcripts demonstrated an attempt to skip intermediate reasoning steps, similar to the unfaithful error recovery behavior observed in the paper.
These findings offer preliminary evidence that the errors introduced in the paper and the observed error recovery behaviors occur naturally.
Appendix C Full error recovery results
C.1 Experiment 1 full results
@@
Figure 10: Faithful recovery rates (as a proportion of all responses) from small and large errors. Error bars indicate 95% binomial confidence intervals.
Figure 11: Unfaithful recovery rates (as a proportion of all responses) from small and large errors. Error bars indicate 95% binomial confidence intervals.
@@
Table 2: Experiment 1 numerical results, performed on GPT-4. All percentages are rounded to 2 decimal places.
Dataset Model Error Amount n Total Error Recovery (%) Faithful Recovery (%) Unfaithful Recovery (%)
Complete Hallucination (%) Partial Hallucination (%)
MultiArith GPT-4 Small Error 247 77.33 72.06 3.64 1.62
Large Error 247 100.00 95.95 0.81 3.24
Claude-3 Opus Small Error 288 67.36 4.86 57.64 4.86
Large Error 288 96.53 3.12 81.94 11.46
Llama-3 70B Small Error 290 45.52 6.55 37.93 1.03
Large Error 290 91.72 17.59 45.52 28.62
ASDiv GPT-4 Small Error 228 75.44 46.93 24.56 3.95
Large Error 228 97.37 85.96 7.89 3.51
Claude-3 Opus Small Error 196 64.29 3.06 53.06 8.16
Large Error 196 86.22 1.53 72.45 12.24
Llama-3 70B Small Error 241 55.19 6.64 47.30 1.24
Large Error 241 91.29 21.58 55.19 14.52
SVAMP GPT-4 Small Error 257 72.76 51.75 19.07 1.95
Large Error 257 93.77 86.38 6.23 1.17
Claude-3 Opus Small Error 272 58.09 1.47 47.06 9.56
Large Error 272 83.09 1.47 70.22 11.40
Llama-3 70B Small Error 284 62.32 2.11 56.69 3.52
Large Error 284 93.31 19.01 50.35 23.94
GSM8K GPT-4 Small Error 229 79.04 50.66 6.11 22.27
Large Error 227 97.80 81.06 6.61 10.13
Claude-3 Opus Small Error 223 67.71 2.69 62.78 2.24
Large Error 223 89.69 1.79 80.72 7.17
Llama-3 70B Small Error 271 31.73 2.58 25.83 3.32
Large Error 271 79.34 7.01 40.96 31.37
C.2 Experiment 2 full results
Figure 12: Overall error recovery rates (as a proportion of all responses) from textual adjustments. Error bars indicate 95% binomial confidence intervals.
Figure 13: Faithful recovery rates (as a proportion of all responses) from textual adjustments. Error bars indicate 95% binomial confidence intervals.
Figure 14: Unfaithful recovery rates (as a proportion of all responses) from textual adjustments. Error bars indicate 95% binomial confidence intervals.
Figure 15: Difference recovery rates between error recovery prompt and baseline conditions, as a proportion of all responses. Negative values indicate recoveries occurred more often in the baseline condition. Error bars are 95% confidence intervals.
Table 3: Experiment 2 numerical results for the MultiArith and ASDiv datasets. All percentages are rounded to 2 decimal places.
Dataset Model Text Adjustment n Total Error Recovery (%) Faithful Recovery (%) Unfaithful Recovery (%)
Complete Hallucination (%) Partial Hallucination (%)
MultiArith GPT-4 Baseline 247 77.73 68.42 4.45 4.86
Context Noise 247 87.45 83.00 3.64 0.81
Error Recovery Prompt 247 81.78 75.71 4.45 1.62
Claude-3 Opus Baseline 282 69.15 2.84 60.28 6.03
Context Noise 282 100.00 0.35 99.65 0.00
Error Recovery Prompt 282 75.53 16.67 57.45 1.42
Llama-3 70B Baseline 289 55.36 5.88 37.72 11.76
Context Noise 289 60.55 6.23 53.29 1.04
Error Recovery Prompt 289 57.79 12.11 21.80 23.88
ASDiv GPT-4 Baseline 226 74.34 51.33 19.03 3.98
Context Noise 226 87.17 64.16 23.01 0.00
Error Recovery Prompt 226 80.97 55.31 24.34 1.33
Claude-3 Opus Baseline 204 70.10 2.45 56.86 10.78
Context Noise 203 99.51 0.49 97.04 1.97
Error Recovery Prompt 204 73.53 14.71 56.37 2.45
Llama-3 70B Baseline 240 70.42 3.75 60.83 5.83
Context Noise 240 77.92 2.92 71.67 3.33
Error Recovery Prompt 240 64.58 10.00 40.83 13.75
Table 4: Experiment 2 numerical results for the SVAMP and GSM8K datasets. All percentages are rounded to 2 decimal places.
Dataset Model Text Adjustment n Total Error Recovery (%) Faithful Recovery (%) Unfaithful Recovery (%)
Complete Hallucination (%) Partial Hallucination (%)
SVAMP GPT-4 Baseline 243 76.54 55.14 19.34 2.06
Context Noise 243 88.89 69.14 19.75 0.00
Error Recovery Prompt 243 84.77 58.44 25.51 0.82
Claude-3 Opus Baseline 269 62.45 3.72 51.67 7.06
Context Noise 269 98.51 0.74 97.77 0.00
Error Recovery Prompt 269 75.09 12.27 55.02 7.81
Llama-3 70B Baseline 281 73.31 6.41 54.45 12.46
Context Noise 281 81.85 2.14 75.44 4.27
Error Recovery Prompt 281 67.62 6.76 40.21 20.64
GSM8K GPT-4 Baseline 224 78.57 51.79 9.82 16.96
Context Noise 224 83.48 69.20 12.50 1.79
Error Recovery Prompt 224 79.91 65.18 5.36 9.38
Claude-3 Opus Baseline 220 69.55 3.18 61.36 5.00
Context Noise 220 98.64 1.36 95.91 1.36
Error Recovery Prompt 220 76.82 15.91 56.82 4.09
Llama-3 70B Baseline 263 42.21 1.14 34.22 6.84
Context Noise 262 52.29 0.76 48.85 2.67
Error Recovery Prompt 263 42.59 4.18 25.86 12.55
C.3 Experiment 3 full results
Figure 16: Total error recovery rates (as a proportion of all responses) for GPT-4 for each error position. Error bars indicate 95% binomial confidence intervals.
Table 5: Experiment 3 numerical results. Due to manual annotation costs, this experiment was only evaluated on GPT-4. The ”Calculation Error” results are the ”Baseline” error recovery results from Experiment 3 (Tables 3 and 4). All percentages are rounded to 2 decimal places.
Dataset Model Error
@@ C.3 Experiment 3 full results
MultiArith GPT-4 Copy Error 295 99.32 72.54 22.37 4.41
Calculation Error 247 77.73 68.42 4.45 4.86
Propagated Calculation Error 295 32.54 25.42 6.44 0.68
ASDiv GPT-4 Copy Error 274 97.45 66.79 20.80 9.85
Calculation Error 226 74.34 51.33 19.03 3.98
Propagated Calculation Error 203 52.71 16.26 35.96 0.49
SVAMP GPT-4 Copy Error 262 97.71 67.56 20.99 9.16
Calculation Error 243 76.54 55.14 19.34 2.06
Propagated Calculation Error 235 54.47 13.62 40.85 0.00
GSM8K GPT-4 Copy Error 279 94.98 58.06 20.79 16.13
Calculation Error 224 78.57 51.79 9.82 16.96
Propagated Calculation Error 242 26.45 12.81 13.22 0.41
Appendix D Sensitivity Analysis